### 01 Jan 2024 v1

@@ 1 Introduction
Compared to those papers, I introduce the possibility of players erasing signals from their records. I show that players’ strategic manipulations of records lead to complete breakdowns in cooperation either when players have long expected lifespans or when their actions are strategic complements, but not when their expected lifespans are intermediate and their actions are substitutes. These conclusions stand in contrast to the ones in the existing literature that higher discount factors and strategic complementarities facilitate cooperation. [^3] Wiseman (2017) and Sandroni and Urgun (2018) show that higher discount factors can undermine cooperation incentives. In contrast to the current paper that focuses on repeated games, their results are obtained in stochastic games with absorbing states. My analysis also highlights the different strategic implications of players’ time preference and their survival probability. By contrast, the two are equivalent in existing models of community enforcement since the equilibrium outcomes depend only on the product of the two, i.e., the player’s discount factor.
Friedman and Resnick (2001) study repeated prisoner’s dilemma without any complementarity or substituability in which players can erase all signals in their records and restart with a fresh record without any signal. In contrast, the players in my model can decide whether to erase each individual signal, which gives them more flexibility in manipulating their records. Compared to their result that cooperation is feasible when players have sufficiently high discount factors, I show that players’ ability to sustain cooperation is not monotone with respect to their discount factors. In an earlier work, Pei (2023) studies a reputation model in which a long-lived player can erase actions from their records. In contrast to the current paper that focuses on games with complete information but allows for imperfect monitoring and multiple long-lived players, that paper studies incomplete information games but with perfect monitoring and only one long-lived player.
This paper is also related to some recent works on dynamic information censoring such as Smirnov and Starkov (2022), Sun (2023), and Hauser (2023). Unlike those papers in which players’ payoffs depend on some unknown state of the world, their payoffs depend only on the action profile in my model.
@@ 2 Model
For players in population $i$, their strategy is given by $\sigma_{i}\equiv(\sigma_{i}^{a},\sigma_{i}^{e})$, where $\sigma_{i}^{a}:\prod_{i=1}^{n}R_{i}\rightarrow\Delta(A_{i})$ is a mapping from his current-period record and the current-period records of other players in his own group to a distribution over his actions and $\sigma_{i}^{e}:R_{i}\times S_{i}\rightarrow[0,1]$ is a mapping from his current-period record and the realized $s_{i}$ in that period to the probability with which he erases that signal realization from his record.
@@
In my baseline model, each player can only erase his signal in period $t$ by the end of period $t$ but cannot do so after period $t$. All my results are robust when players can also erase his period-$t$ signal after period $t$. My baseline model also assumes that players incur no cost to erase signals. My results are robust when they incur an additive cost of erasing signals, as long as that cost is small enough.
@@ 4 Cooperation Between Players with Intermediate Lifespans
My definition of purifiable equilibria in repeated games follows from that in Bhaskar and Thomas (2019). A perturbed stage game is denoted by $\Gamma(\varepsilon)$, where $\varepsilon>0$ is a parameter. Player $i$’s stage-game payoff from playing action $a_{i}\in A_{i}$ is augmented by $\varepsilon z_{i}(a_{i})$ where $z_{i}(a_{i})$ is a random variable with a bounded support. The random variables $z_{i}(a_{i})$ are independently distributed across actions, players, and periods, and these distributions are atomless. Each period, before player $i$ acts, he observes the realized shocks $\{z_{i}(a_{i})\}_{a_{i}\in A_{i}}$ for the current period but not the payoff shocks of the other players and those in future periods. An equilibrium $\sigma$ of the unperturbed repeated game is purifiable if for every sequence $\varepsilon_{n}\rightarrow 0$, there exist a sequence of equilibria $\sigma(\varepsilon_{n})$ of the perturbed repeated games $\Gamma(\varepsilon_{n})$ that converge to $\sigma$.
I start from a result showing that it is impossible to sustain any cooperation in any purifiable equilibrium in the supermodular prisoner’s dilemma.
@@
Suppose $g\leq l$. There is no purifiable equilibrium in which the average probability of playing $C$ is strictly positive regardless of $\{\widehat{\delta}_{i},\overline{\delta}_{i}\}_{i\in\{1,2\}}$ and the monitoring technology.
@@
Second, the comparison between Theorems 2 and 3 suggests that players can sustain cooperation in purifiable equilibria in the submodular prisoner’s dilemma but cannot sustain any cooperation in the supermodular prisoner’s dilemma. [^4] When payoffs are supermodular, one can construct equilibria in which the average probability with which players play $C$ is strictly positive as long as $g\leq l\leq 1+g$ although those equilibria are not purifiable. This stands in contrast to the takeaways in Takahashi (2010), Heller and Mohlin (2018), and Clark, Fudenberg and Wolitzky (2021) that when records are first order, it is easier to sustain cooperation when payoffs are supermodular compared to the case where payoffs are submodular. [^5] Heller and Mohlin (2018) study a community enforcement model in which there are mixed-strategy commitment types and each player can observe $k$ random samples of their partner’s past play before choosing their actions. Takahashi (2010) and Clark, Fudenberg and Wolitzky (2021)’s negative results on the submodular prisoner’s dilemma with first-order records restrict attention to strict equilibria, which is a subset of purifiable equilibria considered in the current paper.

### 14 Mar 2024 v2

@@ 1 Introduction
Compared to those papers, I introduce the possibility of players erasing signals from their records. I show that players’ strategic manipulations of records lead to complete breakdowns in cooperation either when players have long expected lifespans or when their actions are strategic complements, but not when their expected lifespans are intermediate and their actions are substitutes. These conclusions stand in contrast to the ones in the existing literature that higher discount factors and strategic complementarities facilitate cooperation. [^3] Wiseman (2017) and Sandroni and Urgun (2018) show that higher discount factors can undermine cooperation incentives. In contrast to the current paper that focuses on repeated games, their results are obtained in stochastic games with absorbing states. My analysis also highlights the different strategic implications of players’ time preference and their survival probability, which to the best of my knowledge, is novel in the community enforcement literature. By contrast, the two are equivalent in existing models of community enforcement since the equilibrium outcomes depend only on the product of the two, i.e., the player’s discount factor.
Friedman and Resnick (2001) study repeated prisoner’s dilemma without any complementarity or substituability in which players can erase all signals in their records and restart with a fresh record without any signal. [^4] Ghosh and Ray (1996) and Fujiwara-Greve and Okuno-Fujiwara (2009) study repeated games with voluntary separation, in which players can build long-term relationships with their current partners and after a player defects, he will be matched with a new partner from the pool of unmatched players. Their models stand in contrast to mine in which players are being rematched in every period, i.e., there is no long-term relationships between players, and moreover, each player has the flexibility to decide whether to include each new signal in their record. In contrast, the players in my model can decide whether to erase each individual signal, which gives them more flexibility in manipulating their records. Compared to their result that cooperation is feasible when players have sufficiently high discount factors, I show that players’ ability to sustain cooperation is not monotone with respect to their discount factors. In an earlier work, Pei (2023) studies a reputation model in which a long-lived player can erase actions from their records and provides a sharp characterization of players’ equilibrium behaviors and payoffs when the cost of erasing actions is below some cutoff. Although the current paper cannot accommodate incomplete information and does not provide a full characterization of players’ equilibrium behaviors, it allows for multiple long-lived players, a larger class of stage-game payoffs, and imperfect monitoring of players’ actions.
This paper is also related to some recent works on dynamic information censoring such as Smirnov and Starkov (2022), Hauser (2023), and Sun (2024). Unlike those papers in which players’ payoffs depend on some unknown state of the world, their payoffs depend only on the action profile in my model.
@@ 2 Model
For players in population $i$, their strategy is given by $\sigma_{i}\equiv(\sigma_{i}^{a},\sigma_{i}^{e})$, where $\sigma_{i}^{a}:\prod_{i=1}^{n}R_{i}\rightarrow\Delta(A_{i})$ is a mapping from his current-period record and the current-period records of other players in his own group to a distribution over his actions and $\sigma_{i}^{e}:R_{i}\times S_{i}\rightarrow[0,1]$ is a mapping from his current-period record and the realized $s_{i}$ in that period to the probability with which he erases that signal realization from his record. [^5] Although players can observe more information in addition to their records and their current partners’ records (e.g., their previous partners’ records, the signals they erased), focusing on this class of strategies (and equilibria resulting from these strategies) is without loss of generality in the sense that for every equilibrium in which players’ strategies depend on payoff-irrelevant information, there exists an equivalent equilibrium in which all players use strategies that belong to my class and their discounted average payoffs and the average probability with which they take each action remain unchanged.
@@
In my baseline model, each player can only erase his signal in period $t$ by the end of period $t$ but cannot do so after period $t$. All my results are robust when players can also erase his period-$t$ signal after period $t$. My results also extend to the case in which in each period, each player chooses a subset of all the signals that he has generated and shows this subset to his current partners before observing the subset of signals presented by his current-period partners.
My baseline model also assumes that players incur no cost to erase signals. My results are robust when they incur an additive cost of erasing signals, as long as that cost is smaller than some cutoff (which depends only on players’ stage-game payoffs).
@@ 4 Cooperation Between Players with Intermediate Lifespans
My definition of purifiable equilibria in repeated games follows from that in Bhaskar and Thomas (2019). Recall that the stage game in my model is a finite normal-form game $\Gamma\equiv\{I,A,u\}$, which I refer to as the unperturbed stage game. For every $\varepsilon>0$, an $\varepsilon$-perturbed stage game is denoted by $\Gamma(\varepsilon)$, in which the set of players and the set of actions are the same as in the unperturbed stage-game but player $i$’s payoff from playing $a_{i}$ is given by $u_{i}(a_{i},a_{-i})+\varepsilon z_{i}(a_{i})$, where $z_{i}(a_{i})$ is a random variable that is interpreted as a payoff shock that affects player $i$’s payoff from playing $a_{i}$. As in Harsanyi (1973), I assume that the distribution of $z_{i}(a_{i})$ has bounded support and no atom, and that these random variables are independently distributed across actions, players, and periods. Each period, before player $i$ acts, he observes the realizations of his own payoff shocks $\{z_{i}(a_{i})\}_{a_{i}\in A_{i}}$ for the current period but not the realized payoff shocks of the other players and those in future periods. An equilibrium $(\sigma,\mu)$ of the unperturbed repeated game is purifiable if for every sequence $\varepsilon\rightarrow 0$, there exist a sequence of equilibria $(\sigma(\varepsilon),\mu(\varepsilon))$ of the $\varepsilon$-perturbed repeated games $\Gamma(\varepsilon)$ that converge to $(\sigma,\mu)$.
I start from a result which shows that it is impossible to sustain any cooperation in any purifiable equilibrium in the supermodular prisoner’s dilemma under any first-order record. This conclusion can be extended to general monotone-supermodular games where there exists a complete order on each player’s action set under which (i) both players’ payoffs have weakly increasing differences and (ii) each player’s payoff is strictly increasing in their opponent’s action and is strictly decreasing in their own action.
@@
Suppose $g\leq l$. There is no purifiable equilibrium in which the average probability of playing $C$ is strictly positive for any $\{\widehat{\delta}_{i},\overline{\delta}_{i}\}_{i\in\{1,2\}}$ and any first-order record.
@@
Second, the comparison between Theorems 2 and 3 suggests that players can sustain cooperation in purifiable equilibria in the submodular prisoner’s dilemma but cannot sustain any cooperation in the supermodular prisoner’s dilemma. [^6] When payoffs are supermodular, one can construct equilibria in which the average probability with which players play $C$ is strictly positive as long as $g\leq l\leq 1+g$ although those equilibria are not purifiable. This stands in contrast to the takeaways in Takahashi (2010), Heller and Mohlin (2018), and Clark, Fudenberg and Wolitzky (2021) that when records are first order, it is easier to sustain cooperation when payoffs are supermodular compared to the case where payoffs are submodular. [^7] Heller and Mohlin (2018) study a community enforcement model in which there are mixed-strategy commitment types and each player can observe $k$ random samples of their partner’s past play before choosing their actions. Takahashi (2010) and Clark, Fudenberg and Wolitzky (2021)’s negative results on the submodular prisoner’s dilemma with first-order records restrict attention to grim trigger equilibria or strict equilibria, which are subsets of purifiable equilibria considered in the current paper.
@@
In order for the above strategy profile to be part of an equilibrium, players’ expected lifespans cannot be too short. This is because otherwise, their effective discount factors will be too low, which precludes their incentives to cooperate. Their expected lifespans cannot be too long since otherwise, there will be too many seniors in the population who are unwilling to cooperate. Anticipating this, juniors will also have no incentive to cooperate since their continuation value once they become senior is too low.